# An Optimism-based Approach to Online Evaluation of Generative Models

### v1

@@ 5 Online Evaluation based on Fréchet Inception Distance
$\widehat{\Sigma}_{g}^{n}=\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\widehat{\mu}_{g}^{n})(f(x^{i})-\widehat{\mu}_{g}^{n})^{\top},\qquad\widehat{\mu}_{g}^{n}=\frac{1}{n}\sum_{i=1}^{n}f(x^{i})$ (5)
are the empirical covariance matrix and the mean vector, respectively, and $f(x^{i})\in{\mathbb{R}}^{d}$ is the feature of the $i$-th generated image which is extracted by, e.g., the InceptionNet.V3. It has been shown that the empirical FID (4) is biased differently depending on the generator, which makes it difficult to compare different generators for a fixed sample size [9]. Hence, the FID-based evaluation is typically performed for a large batch of generated samples (e.g., $10\sim 50$k) to reduce the effect of the bias, which can be sample-inefficient and costly for a large number of generators.
To enable sample-efficient online evaluation of generative models, we adapt the optimism-based online learning framework to the FID score. To this end, we first derive an optimistic FID score in the following theorem. We defer the theorem’s proof to Appendix A.1.
Theorem 5.1 (Generator-dependent optimistic FID score).
Assume the covariance matrix $\Sigma_{\textup{r}}\succ\bm{0}$ is positive definite. Then, with probability at least $1-\delta$, we have
$\widehat{\textup{FID}}_{g}^{n}=\widetilde{\textup{FID}}_{g}^{n}-{\mathcal{B}}^{n}_{g}\leq\textup{FID}_{g}.$ (6)
Here, the bonus is given by
${\mathcal{B}}_{g}^{n}:=$ $C^{n}_{g}\cdot\sqrt{\frac{2{\mathcal{I}}_{g}}{n}}\left(32\log\left(\frac{24d}{\delta}\right)\right)^{\frac{1}{4}}+\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)}$ (7)
$+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\left(20\sigma_{g}^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}+\frac{2{\mathcal{I}}_{g}}{n}\sqrt{32\log\left(\frac{24d}{\delta}\right)}\right)},$
where $C_{g}^{n}:=2(\|\mu_{g}-\mu_{\textup{r}}\|_{2}+\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2})$, ${\mathcal{I}}_{g}:=\sum_{j=1}^{d}\sum_{l=1}^{d}|\Sigma_{g}[l,j]|$, and $\bm{r}(\Sigma_{g}):=\frac{\textup{Tr}[\Sigma_{g}]}{\|\Sigma_{g}\|_{2}}$.
Algorithm 1 FID-UCB for Generative Model Evaluation
@@
1:  Initialize: $\widehat{\textup{FID}}_{g}\leftarrow-\infty,{\mathcal{H}}_{g}\leftarrow\varnothing,n_{g}\leftarrow 0$ for each $g\in[G]$.
2:  for step $t=1,2,\cdots,T$ do
3:     Pick model $g_{t}\leftarrow\operatorname*{arg\,min}_{g\in{\mathcal{G}}}\widehat{\textup{FID}}_{g}$.
4:     Model $g_{t}$ generates batch samples $\{x_{t}^{j}\in{\mathbb{R}}^{D}\sim g_{t}\}_{j=1}^{b}$.
5:     Update ${\mathcal{H}}_{g_{t}}\leftarrow{\mathcal{H}}_{g_{t}}\cup\{f(x_{t}^{j})\}_{j=1}^{b}$ and visitation $n_{g_{t}}\leftarrow n_{g_{t}}+b$.
6:     Update $\widehat{\mu}_{g_{t}}$ and $\widehat{\Sigma}_{g_{t}}$.
7:     Compute the optimistic FID score
$\widehat{\textup{FID}}_{g_{t}}\leftarrow\|\widehat{\mu}_{g_{t}}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}\left[\widehat{\Sigma}_{g_{t}}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g_{t}}\Sigma_{\textup{r}})^{\frac{1}{2}}\right]-{\mathcal{B}}_{g_{t}}^{n_{g_{t}}}.$
8:  end for
We remark that the bonus term (7) is: 1) generator-dependent through parameters including ${\mathcal{I}}_{g}$, $\sigma_{g}^{2}$, $\textup{Tr}[\Sigma_{g}]$, and $\|\Sigma_{g}\|_{2}$, 2) dimension-free in the sense that the dimension $d$ only appears in the logarithmic term, which facilitates a sample-efficient evaluation. In practice, these parameters can be estimated from the queried data.
@@
Based on Theorem 5.1, we propose FID-UCB in Algorithm 1 as an FID-based online evaluation algorithm. At the beginning, the estimated FID scores are initialized to be negative infinity (line 1), and hence each generator is explored for at least one time. The evaluation proceeds iterative, where at each iteration $t\in[T]$, the evaluator picks model $g_{t}$ with the lowest estimated FID and queries a batch of images from the model (lines $3\sim 4$). Then, the estimated FID of generator $g_{t}$ is updated (lines $5\sim 7$). Next, we show that FID-UCB attains sub-linear regret bound. We defer the theorem’s proof to Appendix A.2.
Theorem 5.2 (Regret of FID-UCB).
Under the same conditions in Theorem A.1, with probability at least $1-\delta$, the regret of the UCB-FID algorithm after $T$ steps is bounded by
$\widetilde{O}\left(\max_{g}\left\{G\cdot{\mathcal{I}}_{g}\cdot\log(T),\sqrt{G({\mathcal{I}}_{g}+\textup{Tr}[\Sigma_{g}])T},(\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\cdot\sigma_{g}\|\Sigma_{g}\|_{2}^{\frac{1}{2}}\bm{r}^{\frac{1}{4}}(\Sigma_{g}))\cdot G^{\frac{1}{4}}T^{\frac{3}{4}}\right\}\right),$ (8)
where logarithmic factors are hidden in the notation $\widetilde{O}(\cdot)$.
6 Online Evaluation based on Inception Score
@@
In this section, we focus on evaluating generative models by Inception score (IS), which is given by $\textup{IS}=\exp[I(X_{g};Y_{g})]$, where $X_{g}\sim p_{g}$ is the generated image, and $Y_{g}$ is the class assigned by the InceptionNet.V3. Given $n\in{\mathbb{N}}_{+}$ generated images $x^{1},\cdots,x^{n}\sim p_{g}$ queried from a generator $g$, the empirical IS is computed by
$\widetilde{\textup{IS}}_{g}^{n}=\exp\left\{\widehat{H}^{n}(Y_{g})-\widehat{H}^{n}(Y_{g}|X_{g})\right\},$ (9)
where
$\widehat{H}^{n}(Y_{g})=H\left(\frac{1}{n}\sum_{i=1}^{n}p_{Y|x^{i}}\right),\ \widehat{H}^{n}(Y_{g}|X_{g})=\frac{1}{n}\sum_{i=1}^{n}H(Y|x^{i}),$ (10)
@@
are the empirical entropy of the marginal $d$-class distribution $p_{Y_{g}}$ and the conditional $d$-class distribution $p_{Y|X_{g}}$ ($d=1000$ for InceptionNet.V3).
@@
$\bm{\epsilon}^{n}_{g}[j]:=\sqrt{\frac{2\widehat{V}^{n}(p_{Y|X_{g}}[j])}{n}\log\left(\frac{4d}{\delta}\right)}+\frac{7}{3(n-1)}\log\left(\frac{4d}{\delta}\right),$ (12)
@@
The optimistic marginal class distribution ensures that ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n}):=-\sum_{j}\widehat{p}_{Y_{g}}^{n}[j]\cdot\log(\widehat{p}_{Y_{g}}^{n}[j])\geq H(Y_{g})$ with high probability (see Lemma B.1 in the Appendix B.3). Next, we derive a generator-dependent optimistic IS in the following theorem. We defer the theorem’s proof to Appendix B.1.
Theorem 6.1 (Generator-dependent Optimistic IS).
@@
$\widehat{\textup{IS}}^{n}_{g}:=\exp\left\{{\mathcal{E}}\left(\widehat{p}_{Y_{g}}^{n}\right)-\widehat{H}^{n}(Y_{g}|X_{g})+\sqrt{\frac{2\widehat{V}^{n}(H(Y_{g}|X_{g}))}{n}\log\left(\frac{4}{\delta}\right)}+\frac{7C^{\prime}}{3(n-1)}\log\left(\frac{4}{\delta}\right)\right\},$ (14)
where $\widehat{H}^{n}(Y_{g}|X_{g})$ is defined in Equation (10), $\widehat{p}_{Y_{g}}^{n}$ the optimistic marginal class distribution (11), parameter $C^{\prime}$ satisfies that $P(|H(Y_{g}|X_{g})|\leq C^{\prime})=1$, and $\widehat{V}^{n}(H(Y_{g}|X_{g}))={\mathbb{V}}(H(Y|x^{1}),\cdots,H(Y|x^{n}))$ is the empirical variance for the entropy of the conditional class distribution. Then, with probability at least $1-\delta$, we have that $\widehat{\textup{IS}}^{n}_{g}\geq\textup{IS}_{g}$.
@@
$\widetilde{p}_{Y_{g_{t}}}\leftarrow\frac{N_{g}\cdot\widetilde{p}_{Y_{g}}+\sum_{j=1}^{b}p_{Y|x_{t}^{i}}}{n_{g_{t}}+b},\ \widehat{H}(Y_{g_{t}}|X_{g_{t}})\leftarrow\frac{N_{g}\cdot\widehat{H}(Y_{g_{t}}|X_{g_{t}})+\sum_{j=1}^{b}H(Y|x_{t}^{i})}{n_{g_{t}}+b}.$
@@
Algorithm 2 IS-UCB for Generative Model Evaluation
Based on Theorem 6.1, we propose IS-UCB in Algorithm 2, an optimism-based online IS evaluation algorithm. In the following theorem, we derive the regret bound of the IS-UCB algorithm, which shows that IS-UCB attains $\widetilde{O}(\sqrt{T})$ regret. We present the theorem’s proof in Appendix B.2.
Theorem 6.2 (Regret of IS-UCB).
$\textup{Regret}(T)\leq\widetilde{O}\left(e^{C}\cdot\max\left\{dm,\left(\frac{d\log d}{4}\right)^{2}\right\}\sqrt{T}+dG\Delta^{-2}\log d\right),$ (15)
where $C:=O(de^{-1}+C^{\prime})$ is the upper bound of $\widehat{\textup{IS}}_{g}^{n}$, $m:=\max_{g\in[G],j\in[d]:p_{Y_{g}}[j]>0}\{|u^{\prime}(p_{Y_{g}}[j])|\}$, $\Delta:=\min_{g\in[G],j\in[d]:p_{Y_{g}}[j]\neq e^{-1}}\{|p_{Y_{g}}[j]-e^{-1}|\}$, and $u(x)=-x\log x$.
@@ 7 Experimental Results
Baseline methods. For both FID-based and IS-based evaluation, we compare the performance of our proposed FID-UCB and IS-UCB with the following two baselines, naive-UCB and the greedy algorithm. Naive-UCB is a simplification of FID-UCB and IS-UCB which is based on data-independent dimension-based concentration bounds, which do not exploit the generated data in evaluating the confidence bound. We present the exact estimators in Appendix D. On the other hand, the greedy algorithm always picks the generator with the lowest empirical FID (4) or the highest empirical IS (9). For all these algorithms, we initialize the estimated FID to be $\infty$- ∞ and the estimated IS to be +$\infty$+ ∞, and hence each generator will be explored at least once.
[…]
@@
Experimental setup and performance metrics. We report three performance metrics at each step $t\in[T]$: 1) total regret, i.e., $\textup{Regret}(t)$, 2) average regret, i.e., $(1/t)\cdot\textup{Regret}(t)$, and 3) optimal pick ratio, i.e., the ratio $(1/t)\cdot\sum_{i=1}^{t}\bm{1}(g_{t}=g^{*})$ of picking the optimal model, which has the lowest empirical FID score or the highest empirical IS score for 50k generated images. For all the experiments, we use a batch size of 5, and the total evaluation step is $T=1,000$. Hence the total generated samples for each trial is $5$k. Results are averaged over 20 trials. Due to the limited space, we postpone parts of the results to Appendix C.
7.1 Empirical Results of Online FID-based Evaluation

### v3

@@ 5 FRÉCHET DISTANCE-BASED ONLINE EVALUATION AND SELECTION
$\widehat{\mu}_{g}^{n}=\frac{1}{n}\sum_{i=1}^{n}f(x^{i}),\ \widehat{\Sigma}_{g}^{n}=\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\widehat{\mu}_{g}^{n})(f(x^{i})-\widehat{\mu}_{g}^{n})^{\top}$
are the mean vector and the empirical covariance matrix, respectively, and $f(x^{i})\in{\mathbb{R}}^{d}$ is the feature of the $i$-th generated image which is extracted by, e.g., the InceptionNet.V3. The FD-based evaluation is typically performed on a large batch of generated samples (10-50 thousands) to reduce the estimation variance, which can be sample-inefficient and costly.
To enable sample-efficient online evaluation of generative models, we adapt the optimism-based online learning framework to the FD score. To this end, we first derive an optimistic FD score in the following theorem. We defer the theorem’s proof to Appendix A.1.
Theorem 1 (Optimistic FD score).
Assume for any generator $g$, the (random) embedding $f(X_{g})\sim{\mathcal{N}}(\mu_{g},\Sigma_{g})$ follows a multivariate Gaussian, and the covariance matrix $\Sigma_{\textup{r}}$ of the real data is positive definite. Then, with probability at least $1-\delta$, we have
$\widehat{\textup{FD}}_{g}^{n}=\widetilde{\textup{FD}}_{g}^{n}-{\mathcal{B}}^{n}_{g}\leq\textup{FD}_{g}$ (5)
for any $n\geq 4\bm{r}(\Sigma_{g})+\log(3/\delta)$, where $\bm{r}(\Sigma_{g}):=\frac{\textup{Tr}[\Sigma_{g}]}{\|\Sigma_{g}\|_{2}}$ is the effective rank of the covariance matrix $\Sigma_{g}$. In addition, the bonus is given by
${\mathcal{B}}_{g}^{n}:=$ $2\Delta_{\mu_{g}}^{n}\cdot\left(\Delta_{\mu_{g}}^{n}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}\right)+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\Delta_{\Sigma_{g}}^{n}}$ (6)
$+\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6}{\delta}\right)}+\frac{8\|\Sigma_{g}\|_{2}}{n}\log\left(\frac{6}{\delta}\right),$
where $\Delta_{\mu_{g}}^{n}$ and $\Delta_{\Sigma_{g}}^{n}$ are defined in (23) and (24), respectively.
Avg. Regret
Figure 3: Online IS-based evaluation and selection among standard generative models: The $x$-axis is the number of steps. At each step, the algorithm samples a batch of five generated images from the chosen model. Results are averaged over 20 trials.
@@
Remark 1 (Model-dependent parameters).
The bonus term (6) involves model-dependent parameters $\textup{Tr}[\Sigma_{g}]$ and $\|\Sigma_{g}\|_{2}$ of the covariance matrix. Note that the presence of the norms of $\Sigma_{g}$ in the concentration bound for FD is inevitable, because, without any assumptions on the norm of $\Sigma_{g}$, the concentration bound cannot have a finite value. To address this, two approaches can be considered: 1) when the norm of the data embeddings is bounded, e.g., the normalized CLIP embeddings where $\|X_{g}\|_{2}\leq 1$, we can substitute $\textup{Tr}[\Sigma_{g}]$ and $\|\Sigma_{g}\|_{2}$ in the above bound with the embedding upper-bound norm. In this case, we can extend our analysis beyond multivariate Gaussians (Theorem 5 in the Appendix). 2) On the other hand, when using the InceptionV3 and DINOv2 embeddings that have generally unbounded output norm, we can use the already generated data to estimate the parameters in the upper-bound. This approach avoids any offline estimation of the parameters required to have an upper-bound and preserves the online format of the FD-UCB algorithm, and conditioned that the estimated norms converge fast to the underlying value, will provide a correct solution. Our numerical results indicate that with only 50 samples, the norm terms in the bound will be relatively close to their actual values. See Part 3 in Appendix C for our experimental validation.
Algorithm 1 FD-UCB
0:  set of models ${\mathcal{G}}$, image embedding $f$, step $T\in{\mathbb{N}}_{+}$, $\mu_{\textup{r}}\in{\mathbb{R}}^{d}$ and $\Sigma_{\textup{r}}\in{\mathbb{S}}^{d}$ of real data, failure probability $\delta\leftarrow\delta/T$, batch size $b\in{\mathbb{N}}_{+}$, number of burn-in samples $N\geq\max_{g\in{\mathcal{G}}}\{4{\bm{r}}(\Sigma_{g})+\log(3/\delta)\}$.
@@
1:  Initialize: $\widehat{\textup{FD}}_{g}\leftarrow-\infty,{\mathcal{H}}_{g}\leftarrow\varnothing,n_{g}\leftarrow 0$ for each $g\in{\mathcal{G}}$.
2:  for model $g\in{\mathcal{G}}$ do
3:     Generate $M$ samples $\{x_{0}^{j}\sim p_{g}\}_{j=1}^{N}$.
4:     Update ${\mathcal{H}}_{g}\leftarrow{\mathcal{H}}_{g}\cup\{f(x_{0}^{j})\}_{j=1}^{N}$ and visitation $n_{g}\leftarrow N$.
5:  end for
6:  for step $t=1,2,\cdots,T$ do
7:     Pick model $g_{t}\leftarrow\operatorname*{arg\,min}_{g\in{\mathcal{G}}}\widehat{\textup{FD}}_{g}$.
8:     Model $g_{t}$ generates batch samples $\{x_{t}^{j}\sim g_{t}\}_{j=1}^{b}$.
9:     Update ${\mathcal{H}}_{g_{t}}\leftarrow{\mathcal{H}}_{g_{t}}\cup\{f(x_{t}^{j})\}_{j=1}^{b}$ and visitation $n_{g_{t}}\leftarrow n_{g_{t}}+b$.
10:     Update $\widehat{\mu}_{g_{t}}$ and $\widehat{\Sigma}_{g_{t}}$.
11:     Compute the optimistic FD score
$\widehat{\textup{FD}}_{g_{t}}\leftarrow$ $-{\mathcal{B}}_{g_{t}}^{n_{g_{t}}}+\|\widehat{\mu}_{g_{t}}-\mu_{\textup{r}}\|_{2}^{2}$
$+\textup{Tr}\left[\widehat{\Sigma}_{g_{t}}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g_{t}}\Sigma_{\textup{r}})^{\frac{1}{2}}\right],$
where ${\mathcal{B}}^{n}_{gt}$ is given by Equation (6).
12:  end for
@@
Based on Theorem 1, we propose FD-UCB in Algorithm 1 as an FD-based online evaluation algorithm. At the beginning, FD-UCB samples $N$ generated data from each model, after which the algorithm can compute an optimist FD score (5) for each model (lines $2$-$5$). The evaluation proceeds iteratively, where at each iteration $t\in[T]$, the evaluator picks model $g_{t}$ with the lowest estimated FD and queries a batch of images from the model (lines $7$-$8$). Then, the estimated FD of generator $g_{t}$ is updated (line $11$). It can be shown that FD-UCB attains sub-linear regret bound, which is formalized in Theorem 3 in Appendix A.2.
6 INCEPTION SCORE-BASED ONLINE EVALUATION AND SELECTION
@@
In this section, we focus on evaluating generative models by Inception score (IS), which is given by $\textup{IS}=\exp[I(X_{g};Y_{g})]$, where $X_{g}\sim p_{g}$ is the generated image, and $Y_{g}$ is the class assigned by the InceptionNet.V3. Given $n\in{\mathbb{N}}_{+}$ generated images $x^{1},\cdots,x^{n}\sim p_{g}$ queried from a generator $g$, we denote by
$\widehat{H}^{n}(Y_{g})=H\left(\frac{1}{n}\sum_{i=1}^{n}p_{Y|x^{i}}\right),$ (7)
$\widehat{H}^{n}(Y_{g}|X_{g})=\frac{1}{n}\sum_{i=1}^{n}H(Y|x^{i}),$ (8)
@@
the empirical entropy of the marginal $d$-class distribution $p_{Y_{g}}$ and the conditional $d$-class distribution $p_{Y|X_{g}}$, respectively. Then, the empirical IS is computed by
$\widetilde{\textup{IS}}_{g}^{n}=\exp\left\{\widehat{H}^{n}(Y_{g})-\widehat{H}^{n}(Y_{g}|X_{g})\right\}.$ (9)
@@
$\bm{\epsilon}^{n}_{g}[j]:=$ $\sqrt{\frac{2\widehat{V}^{n}(p_{Y|X_{g}}[j])}{n}\log\left(\frac{4d}{\delta}\right)}$ (11)
$+\frac{7}{3(n-1)}\log\left(\frac{4d}{\delta}\right),$
@@
Particularly, the optimistic marginal class distribution ensures that ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n})\geq H(Y_{g})$ with high probability (see Lemma 11 in the Appendix), where we define
${\mathcal{E}}(z):=-\sum_{j}z[j]\cdot\log(z[j])$
for any $d$-dimensional vector $z\succ\bm{0}$. Next, we derive a generator-dependent optimistic IS in the following theorem. We defer the theorem’s proof to Appendix B.1.
Theorem 2 (Optimistic IS).
@@
[…]
$+\sqrt{\frac{2\widehat{V}^{n}(H(Y_{g}|X_{g}))}{n}\log\left(\frac{4d}{\delta}\right)}$
$+\frac{7\log d}{3(n-1)}\log\left(\frac{4d}{\delta}\right)\Bigg{\}},$
where $\widehat{H}^{n}(Y_{g}|X_{g})$ is defined in Equation (8), $\widehat{p}_{Y_{g}}^{n}$ the optimistic marginal class distribution (10), and $\widehat{V}^{n}(H(Y_{g}|X_{g}))={\mathbb{V}}(H(Y|x^{1}),\cdots,H(Y|x^{n}))$ is the empirical variance for the entropy of the conditional class distribution. Then, with probability at least $1-\delta$, we have that $\widehat{\textup{IS}}^{n}_{g}\geq\textup{IS}_{g}$.
Figure 4: Online IS-based evaluation and selection among variance-controlled FFHQ models: IS-UCB can identify models that generate images with more diversity. Results are averaged over 20 trials.
[…]
@@
$\widetilde{p}_{Y_{g_{t}}}\leftarrow\frac{N_{g}\cdot\widetilde{p}_{Y_{g}}+\sum_{j=1}^{b}p_{Y|x_{t}^{i}}}{n_{g_{t}}+b},$
$\widehat{H}(Y_{g_{t}}|X_{g_{t}})\leftarrow\frac{N_{g}\cdot\widehat{H}(Y_{g_{t}}|X_{g_{t}})+\sum_{j=1}^{b}H(Y|x_{t}^{i})}{n_{g_{t}}+b}.$
@@
Algorithm 2 IS-UCB
Based on Theorem 2, we propose IS-UCB in Algorithm 2, an optimism-based online IS evaluation algorithm. Due to limited space in the main text, we derive the regret bound of the IS-UCB algorithm in Appendix B.2, which shows that IS-UCB attains $\widetilde{O}(\sqrt{T})$ regret.
@@ 7 EXPERIMENTAL RESULTS
In this section, we present numerical results for the FD-UCB and IS-UCB algorithms. Our extensive experiments across standard image datasets, generative modeling frameworks, and image data embeddings demonstrate the effectiveness of the proposed online learning-based approaches to model evaluation and selection. Experimental details and additional results can be found in Appendix C.
@@
Baselines. For both FD-based and IS-based evaluation, we compare the performance of our proposed FD-UCB and IS-UCB with three baselines: Naive-UCB, Greedy algorithm, and Random algorithm. Naive-UCB is a simplification of FD-UCB and IS-UCB which replaces the generator-dependent variables in the bonus function with data-independent and dimension-based terms, which does not exploit the generated data in evaluating the confidence bound. On the other hand, the Greedy algorithm always picks the generator with the lowest empirical FD (4) or the highest empirical IS (9), and the Random algorithm selects the model randomly. For a fair comparison, we set the burn-in samples of FD-UCB to be zero. Additionally, for all these algorithms, each generator will be explored once at the beginning.
@@
Experimental setup and performance metrics. We report two performance metrics at each step $t\in[T]$: 1) average regret (Avg. Regret), i.e., $(1/t)\cdot\textup{Regret}(t)$, and 2) optimal pick ratio (OPR), i.e., the ratio $(1/t)\cdot\sum_{i=1}^{t}\bm{1}(g_{t}=g^{*})$ of picking the optimal model, which has the lowest empirical FD score or the highest empirical IS score for 50k generated images. For all the experiments, we use a batch size of 5, and the total evaluation step is $T=1,000$. Hence the total generated samples for each trial is $5$k. Results are averaged over 20 trials.
Datasets and generators. We evaluate the above algorithms on standard image datasets, including CIFAR10 (Krizhevsky et al.,, 2009), ImageNet (Deng et al.,, 2009), FFHQ (Kazemi and Sullivan,, 2014), and AFHQ (Choi et al.,, 2020). For the first three datasets, we consider evaluation and selection among standard generative models, such as BigGAN-Deep (Brock et al.,, 2019), iDDPM-DDIM (Nichol and Dhariwal,, 2021), and Efficient-vdVAE (Hazami et al.,, 2022). Details of the chosen models can be found in Appendix C. Additionally, we consider variance-controlled generators for the FFHQ and AFHQ datasets, where we apply the standard truncation technique (Kynkäänniemi et al.,, 2019) to the pretrained StyleGAN2-ADA model (Karras et al.,, 2020) [^1] The repository can be found at github.com/NVlabs/stylegan2-ada-pytorch and synthesize $G=20$ models. Each model generates images from truncated random noises with parameters vary from $0.01$ to $0.1$. A small (large) truncation parameter can lead to generated samples with low (high) diversity but high (low) quality.
7.1 Results of Online FD-based Evaluation and Selection
FID-based evaluation and selection. We first report results for the setup in Figure 1, where we select among three standard generative models on the CIFAR10 dataset, including LOGAN (Wu et al.,, 2020), RESFLOW (Chen et al.,, 2019), and iDDPM-DDiM (Nichol and Dhariwal,, 2021). Our proposed FD-UCB can quickly identify the best model with fewer queries to the suboptimal models (Figure 5).
Figure 5: Online FD-based evaluation and selection among three CIFAR10 models, including LOGAN, RESFLOW, and iDDPM-DDiM (Figure 1). The image data embeddings are extracted by InceptionV3.Net. Results are averaged over 20 trials.